3. (20 points) Suppose f(2) = -3, and f(2) = 4.de Show Dollar svigt Joutro a. Let g(t) = (x2+x)f(x). First find g(x), then
Please do all parts(so a-d) Thank you.
Answer #1

Solutions- (2)=-3 and I(2-4 ts 24-15 f(2)=-3 (a) goo - (x*+w7fczes diffierentiate geoes with respect to X3- gase (x2+x) plAt X-23- 912) = f(2) Sin (%) + fila) los ( W) (-2, gla). 4**. 31 gla) + 31 x 13 2.68 (d, genere comunaling dax da d(x2+x9 f

Please thumbs up if it was helpful will be glad to know;)